Paris:

What is the midpoint of (2,-2, 4) and (-2,0, 6)?

Paris:

@Vocaloid @JustSaiyan @Shadow @dude

Shadow:

\[M = ( \frac{ x _{1} + x _{2}}{ 2 } , \frac{ y _{1} + y _{2} }{ 2 })\]

Shadow:

Is this third dimension

Paris:

Not sure what that means

Shadow:

3D

Paris:

Oh yeah sorry I read that wrong

Shadow:

\[M = ( \frac{ x _{1} + x _{2}}{ 2 } , \frac{ y _{1} + y _{2} }{ 2 }, \frac{ z_{1} + z_{2} }{ 2})\]

Shadow:

Use that one instead

Paris:

So the is the layout of it?

Paris:

Okay thanks you!!

Shadow:

The formula for the midpoint of coordinates in 3D

Paris:

Thank you so much!

Shadow:

No problem

Paris:

@Shadow did I type it right?

1 attachment
Shadow:

It's not supposed to be to the first and second power \[x_{2} \neq x^{2}\]

Shadow:

Those numbers on the bottom just indicate that it's the first or second x, y, or z coordinate.

Paris:

Oh okay

Shadow:

Besides that it looks good, you can add then divide.

Paris:

I'm using this calculator and it keeps saying they I'm missing something I'm really confused.

Paris:

Never mind @Shadow I got it thank you a bunch for helping!!

Shadow:

No problem